OpenStudy (anonymous):

Angle A has a measure of 42 degrees. If Angle B and Angle A are complementary, what is the measure of Angle B?

OpenStudy (anonymous):

if two angles are complementary then it means they add up to 90 degrees.... so in your case a+b=90 deg and a=42 deg, so b=90-42 = 48 degrees.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np dude